OpenStudy (anonymous):

(Logarithms) solve for x : \[\x^{3/4(log_{2}x)^{2}+log2(x)−5/4}=sqrt{2}\]

OpenStudy (anonymous):

@Wired help!!

OpenStudy (anonymous):

take log both sides..and simplfy

OpenStudy (anonymous):

and put logx=y to make a quad equation

OpenStudy (anonymous):

find y and take its log inverse to get x

OpenStudy (anonymous):

i was doing the same, But i am stuck at a step.

OpenStudy (anonymous):

it will be a cubic eq

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am getting a cubic eq.

OpenStudy (anonymous):

i am not able to solve that cubic eq.

OpenStudy (anonymous):

@himanshu31 what eq. have you got?

OpenStudy (anonymous):

is it 3y3+4y2-g5-2log2=0

OpenStudy (anonymous):

thats 5y

OpenStudy (anonymous):

i got 3t^3 +4t^2 - 5t - 2 = 0

OpenStudy (anonymous):

when u take log both side the right side will 1/2(log2)

OpenStudy (zzr0ck3r):

hit enter wired

OpenStudy (anonymous):

@himanshu31 what have you taken for y??

OpenStudy (anonymous):

y=logx

OpenStudy (anonymous):

how do u get '2'?

OpenStudy (anonymous):

i have taken t=log(base2)x

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

On my RHS i have 1/log(base2)x

OpenStudy (anonymous):

is it sqrt2 or sqrtx in question

OpenStudy (anonymous):

sqrt2

OpenStudy (anonymous):

@zzr0ck3r Man, I haven't done logarithms in years, give me a min lol! In the middle of the exponent is that: \[\log2{x} \] or \[\log_2{x}\]

OpenStudy (zzr0ck3r):

im just playin man:)

OpenStudy (anonymous):

LOL I know :)

OpenStudy (anonymous):

log(base2)x @Wired

OpenStudy (anonymous):

ur cubic eq is wrong ..@shubham.bagrecha

OpenStudy (anonymous):

@himanshu31 i got this as one of my steps. \[3(\log_{2}x) ^{2}+\log_{2}x^{2}-5/2=\log_{x}2 \]

OpenStudy (zzr0ck3r):

im to lazy to grab a pencil and join in

OpenStudy (anonymous):

how log(base2)x^2 and how log(base x)2..???

OpenStudy (anonymous):

for removing root, squared both sides.

OpenStudy (anonymous):

when u take log both side the root will become 1/2 .. there is no need to square both sides

OpenStudy (anonymous):

i have not taken log both side. I just used the fact that \[a^{x}=b, x=\log_{a}b\]

OpenStudy (anonymous):

@Wired Have you got something?

OpenStudy (anonymous):

Still running through it. Like I said, it's been a while lol.

OpenStudy (anonymous):

\[x ^{(3/4)(\log_{2}x)^{2} + \log_2{x}-(5/4)} = \sqrt{2}\] \[[(3/4)(\log_{2}x)^{2} + \log_2{x}-(5/4)]log_2{x} = log_2{\sqrt{2}}\] \[4[(3/4)(\log_{2}x)^{2} + \log_2{x}-(5/4)]log_2{x} = 4log_2{\sqrt{2}}\] \[[3(\log_{2}x)^{2} + 4\log_2{x}-5]log_2{x} = 4log_2{\sqrt{2}}\] \[Y=log_{2}x\] \[(3Y^{2} + 4Y-5)Y = 4log_2{\sqrt{2}}\] \[3Y^{3} + 4Y^{2}-5Y - 4log_2{\sqrt{2}} = 0\] \[3Y^{3} + 4Y^{2}-5Y - 4(0.5) = 0\] \[3Y^{3} + 4Y^{2}-5Y - 2 = 0\]

OpenStudy (anonymous):

hey @Wired i want the value of Y me and himanshu also got the same eq. as that of you.

OpenStudy (shubhamsrg):

3y^3 -3y^2 +7y^2 +2y -7y -2 =0 3y^2(y-1) + 7y(y-1) + 2(y-1)=0 try now..

OpenStudy (anonymous):

@shubhamsrg great!!!!! Is there any method to solve that or just trial and error??

OpenStudy (shubhamsrg):

i used hit and trial lol!! :P

OpenStudy (shubhamsrg):

1 was easy to check in this case though

OpenStudy (anonymous):

Trial and error will kill you in a test though. There's a formula / trick, trying to find it.

OpenStudy (shubhamsrg):

you may use rational root theorem also (does not work everytime)

OpenStudy (anonymous):

There's just too many formulas for it. Some are insanely huge. Skipping that :) WolframAlpha says Y = -2, -(1/3), 1 \[Y = \log_2{x}\] \[\log_2{x} = -2\] \[x=2^{-2}=\frac{1}{2^{2}} = \frac{1}{4}\] \[\log_2{x} = -\frac{1}{3}\] \[x=2^{\frac{-1}{3}}=\frac{1}{2^{\frac{1}{3}}}=\frac{1}{\sqrt[3]{2}}\] \[\log_2{x} = 1\] \[x=2^{1}=2\]